# NeoBERT: A Next-Generation BERT

### v1

@@ 4 Effect of Design Choices
Figure 1: GLUE ablation scores on the development set. Modifications in grey are not included in the subsequent models. Increasing data size and diversity leads to the highest relative improvement ($M2$, $+3.6\%$), followed by the model size ($M7$, $+2.9\%$). Packing the sequences and using the LLaMA 2 tokenizer cause the largest relative drops ($M6$, $-2.9\%$, $M3$, $-2.1\%$).
@@
In $M3$, replacing the Google WordPiece tokenizer with LLaMA BPE results in a $-2.1\%$ performance decrease. We hypothesize that while the heterogeneous and multilingual nature of the LLaMA 2 vocabulary enhances broader applicability in decoders, it trades off performance for more compact encoder representations. In $M6$, we un-pad the sequences by concatenating samples of the same batch. While this removes unnecessary computation on padding tokens, packing sequences without accounting for cross-sequence attention results in a relative performance drop of $-2.8\%$. Although this modification was discarded from our subsequent ablations, we incorporate methods of un-padding with accurate cross-attention in our released codebase, following Kundu et al. (2024).
@@ Results
Table 4: MTEB scores on the English subset after 2,000 steps of fine-tuning with contrastive learning.
@@ 5.3 Sequence Length
Although NeoBERT1024 was trained exclusively on sequences of up to $1,024$ tokens, it generalizes effectively to context lengths approaching 3,000 tokens. This demonstrates the robustness of RoPE embeddings to out-of-distribution inputs. Moreover, after an additional 50k training steps with sequences up to $4,096$ tokens, NeoBERT4096 successfully models longer sequences. This approach provides a compute-efficient strategy for extending the model’s maximum context window beyond its original length.
@@ 7 Conclusion
We introduced NeoBERT, a state-of-the-art encoder pre-trained from scratch with the latest advancements in language modeling, architecture, and data selection. To ensure rigorous validation, we systematically evaluated every design choice by fully training and benchmarking ten distinct models in controlled settings. On GLUE, NeoBERT outperforms BERTlarge and NomicBERT and is comparable with RoBERTalarge despite being 100M parameters smaller and supporting sequences eight times longer. To further validate its effectiveness, we conducted a comprehensive evaluation on MTEB, carefully isolating the effects of pre-training and fine-tuning to provide a fair comparison against the best open-source embedding models. Under identical fine-tuning conditions, NeoBERT consistently outperforms all baselines. With its unparalleled efficiency, optimal depth-to-width, and plug-and-play compatibility, NeoBERT represents the next generation of encoder models. To foster transparency and collaboration, we release all code, data, model checkpoints, and training scripts, making NeoBERT the only fully open-source model of its kind.
@@ Appendix C GLUE
Following BERT, we exclude WNLI from our evaluation [^3] See 12 in https://gluebenchmark.com/faq. For tasks with two scores and for MNLI matched and mismatched, we report the average between the two metrics.
@@ D.3 Task instructions
Appendix E Efficiency

### v2

@@ 4 Effect of Design Choices
Figure 1: GLUE ablation scores on the development set. All modifications are cumulative, except for $M3$ and $M6$ in grey, which are not included in the subsequent models. Increasing data size and diversity leads to the highest relative improvement ($M2$, $+3.6\%$), followed by the model size ($M7$, $+2.9\%$). Packing the sequences and using the LLaMA 2 tokenizer cause the largest relative drops ($M6$, $-2.9\%$, $M3$, $-2.1\%$).
@@
In $M3$, replacing the Google WordPiece tokenizer with LLaMA BPE results in a $-2.1\%$ performance decrease. The performance gap observed likely stems from inherent differences between the algorithms, training corpus, and design strategies employed by the two tokenizers. We further discuss this in Appendix B. In $M6$, we unpad the sequences by concatenating samples of the same batch. While this removes unnecessary computation on padding tokens, packing sequences without accounting for cross-sequence attention results in a relative performance drop of $-2.8\%$. Although this modification was discarded from our subsequent ablations, we incorporate methods of un-padding with accurate cross-attention in our released codebase, following Kundu et al. (2024).
@@ Results
Table 4: MTEB(eng, v1) scores after 2,000 steps of fine-tuning with our unified contrastive learning.
@@
12 tasks 11 tasks 3 tasks 4 tasks 15 tasks 10 tasks 1 tasks
@@
5.2.2 Contextual Document Embeddings
While a unified fine-tuning framework enables fair model comparisons, computationally intensive fine-tuning techniques are necessary to achieve top-ranking MTEB scores. We apply one such strategy, Contextual Document Embeddings (CDE)  (Morris & Rush, 2024), to NeoBERT. CDE previously led the MTEB leaderboard with NomicBERT (cde-v1) and ModernBERT (cde-v2) for English models under 400M parameters. NeoBERT outperforms both, achieving an average score of 66.60, the highest in its parameter class. See Appendix G for more details on the fine-tuning.
Table 5: MTEB(eng, v1) scores after fine-tuning with CDE. Ranks on MTEB(eng, v1) for models under 400M parameters as of April 2025.
Model Class. Clust. PairClass. Rerank. Retriev. STS Summ. Avg. Rank
12 tasks 11 tasks 3 tasks 4 tasks 15 tasks 10 tasks 1 tasks
NomicBERT 81.72 48.32 84.69 56.75 53.27 81.64 31.23 65.00 3
ModernBERT 80.62 49.48 85.23 56.94 54.19 83.30 31.60 65.68 2
NeoBERT 82.14 50.29 86.71 57.71 56.37 82.30 30.98 66.60 1
@@ 5.3 Sequence Length
Although NeoBERT1024 was trained exclusively on sequences of up to $1,024$ tokens, it generalizes effectively to context lengths approaching 3,000 tokens. This demonstrates the robustness of RoPE embeddings to out-of-distribution inputs. Moreover, after an additional 50k training steps with sequences up to $4,096$ tokens, NeoBERT4096 successfully models longer sequences. This approach provides a compute-efficient strategy for extending the model’s maximum context window beyond its original length. To further test NeoBERT’s generalization capabilities, we extend the context beyond what was seen during training and present results in Appendix E.
@@ 7 Conclusion
We introduced NeoBERT, a state-of-the-art encoder pre-trained from scratch with the latest advancements in language modeling, architecture, and data selection. To ensure rigorous validation, we systematically evaluated every design choice by fully training and benchmarking ten distinct models in controlled settings. On GLUE, NeoBERT outperforms BERTlarge and NomicBERT and is comparable with RoBERTalarge despite being 100M parameters smaller and supporting sequences eight times longer. To further validate its effectiveness, we conducted a comprehensive evaluation on MTEB, carefully isolating the effects of pre-training and fine-tuning to provide a fair comparison against the best open-source embedding models. Under identical fine-tuning conditions, NeoBERT consistently outperforms all baselines. In addition, when fine-tuned with more advanced techniques, NeoBERT ranks at the top of the MTEB leaderboard for models under 400M parameters. With its unparalleled efficiency, optimal depth-to-width, and plug-and-play compatibility, NeoBERT represents the next generation of encoder models. To foster transparency and collaboration, we release all code, data, model checkpoints, and training scripts, making NeoBERT the only fully open-source model of its kind.
@@ B.1 Baseline
B.2 Tokenizers
We refer to the Llama Wordpiece tokenizer as llama-tok and the BERT BPE tokenizer as bert-tok. Under identical conditions between ablations $M2$ and $M3$, the use of llama-tok over bert-tok decreases the GLUE score by a relative drop of $2.9\%$. Although further investigation would be required to analyze this drop, we highlight several key differences between the two tokenizers.
While llama-tok uses Byte-Pair Encoding (BPE), bert-tok relies on the more expressive WordPiece algorithm, which better models token likelihoods in text. Moreover, their training data largely differs. llama-tok was trained on a dataset mostly composed of the CommonCrawl, whereas bert-tok used the higher-quality BookCorpus and WikiText datasets. Finally, llama-tok is cased and handles unknown UTF-8 characters at the byte level, while bert-tok is uncased and character-level. This leads llama-tok to include redundant tokens (e.g., “The” and “the”) and rare byte sequences, limiting space for more frequent tokens under a finite capacity.
@@ D.3 Task instructions
Appendix E Context Length
We compute the Pseudo-Perplexity on sequences sampled from English Wikipedia. Although NeoBERT has only been exposed to sequences up to 4,096 tokens during training, it correctly models sequences up to 6,000 tokens. Additional training or RoPE extension techniques, such as YaRN (Peng et al., 2023), would be needed to further extend the context window.
Figure 6: Pseudo-Perplexity in function of the sequence length for NeoBERT4096 for extension beyond sequence lengths seen in training. NeoBERT generalizes natively for sequences under 6,000 tokens.
Appendix F Efficiency
@@
Appendix G Contextual Document Embeddings
One popular application for BERT-style models is to generate embedding vectors for use in tasks such as classification, clustering, and retrieval. State-of-the-art embedding models below 400M parameters are trained using contextual embedding techniques (Morris & Rush, 2024). The highest-performing similar-sized embedding models were trained with backbones initialized with NomicBERT (Nussbaum et al., 2024) and more recently ModernBERT (Warner et al., 2024). In this section, we question whether contextual document embeddings (CDE) might perform better when initialized from NeoBERT.
G.1 Method
We train our embedding models using both the contextual batching strategy and contextual architecture described in Morris & Rush (2024).
Contextual batching.
To train a contextual embedding model, we have to first group documents from a larger corpus into miniature ‘pseudo-contexts’ which comprise the batches for training. We do so by minimizing the following batch-clustering objective:
$\min_{\begin{subarray}{c}({\cal B}^{1},\ldots{\cal B}^{B})\\ (c^{1},\ldots,c^{B})\end{subarray}}\sum_{b}\sum_{\begin{subarray}{c}(d,q)\in{\cal B}^{b}\end{subarray}}M((d,q),c^{b})$
Here, $\mathcal{B}_{1}...\mathcal{B}^{B}$ represent the orderings of batches of training data and $c^{*}$ are the corresponding centroids; $M(d,q)$ indicates the distance (typically Euclidean) between embeddings of document $d$ and query $q$ from batch $\mathcal{B}^{b}$.
We can approximate this efficiently using a modified K-Means solver. We follow the same procedure as Morris & Rush (2024) to cluster documents for each training set.
Contextual architecture.
Contextual embedding models in effect use two model backbones: one to embed similar documents in context and another to embed documents conditioned on contextual information. We can express contextual embedding vector $\phi$ using both models $M_{1}$ and $M_{2}$:
$\phi(x;\mathcal{D})=M_{2}(M_{1}(d^{1}),\ldots,M_{1}(d^{J}),E(x_{1}),\ldots,E(x_{T}))$
where $x$ is a document or query comprised of tokens $x_{1}...x_{T}$ and $\mathcal{D}$ is a set of documents from the surrounding context. We train $\phi$ end-to-end by backpropagating through both $M_{1}$ and $M_{2}$ using the custom two-stage gradient-caching procedure described in Morris & Rush (2024).
G.2 Data and hyperparameters
Training stages and data.
We train our models in two stages. The first stage is a large, noisy “contrastive pre-training” stage with $235M$ query-to-document pairs per epoch. The second is a short fine-tuning step on 1.5M high-quality datapoints. We train each stage for three epochs with the Adam optimizer with $1000$ steps of warmup to a learning rate of $55e-$ 5 and linear decay to 00 throughout training.
Instead of mining hard negatives, we cluster the datasets for contextual batching. We embed for clustering and filter hard negatives within a batch using GTE-large (Li et al., 2023a). We set both the batch size and cluster size to $512$. Unlike many contrastive training setups, we neglect to share negatives between GPUs.
We evaluate our contextual embedding models on the $56$ tasks of the English version of MTEB (Muennighoff et al., 2023). To maintain consistency with prior iterations of CDE, we train these models with $512$ in-context documents and a per-document sequence length of $512$; this produces a second-stage input size of $1024$ tokens.
G.3 Results
Our contextual embedding results are detailed in Table 5, averaged across tasks in MTEB. With a score of $66.60$, NeoBERT significantly outperforms ModernBERT in embedding-related tasks; in our case, the gap between NeoBERT and ModernBERT (about $1$ MTEB point) is notably larger than the gap between ModernBERT and NomicBERT (approximately $0.6$).
NeoBERT’s largest improvements are shown in retrieval, reranking, clustering, and classification tasks. We see a small drop in performance in STS compared to ModernBERT, which we hypothesize may relate to the mix of pretraining data.
When compared to other models in MTEB, NeoBERT fine-tuned with CDE is state-of-the-art in its parameter class, outperforming all other similar-sized embedding models as of April 2025, such as GTE with $64.36$ (Li et al., 2023a), BGE with $63.99$ (Chen et al., 2024), and GIST with $64.13$ (Solatorio, 2024).